# ODIN: A Single Model for 2D and 3D Perception

### v1

@@ 4.4 Ablations and Variants
Joint 2D-3D training helps 3D perception We compare joint training of ODIN on sensor RGB-D point clouds from ScanNet and 2D RGB images from COCO to variants trained independently on 2D and 3D data, all initialized from pre-trained COCO weights. Since there are different classes in ScanNet and COCO, we use our open-vocabulary semantic class-decoding head instead of the vanilla closed-vocabulary head. Results in Tab. 4 show that joint training yields a 2.6% absolute improvement in 3D, and causes a similar drop in 2D. We show qualitative results in Fig. 3. The COCO dataset is orders of magnitude larger than ScanNet, and we observe that the model fits faster on ScanNet. As a result, the model keeps improving on COCO while on ScanNet it begins to overfit. This highlights the need to explore smarter strategies for balancing 2D and 3D data during training to prevent overfitting in one domain and underfitting in the other – we leave this for future work. Nevertheless, this experiment highlights the benefits of joint training with 2D datasets for 3D segmentation in ODIN. Note that we do not jointly train on 2D and 3D datasets in any of our other experiments due to computational constraints.
@@
Cross-View fusion is crucial for instance segmentation but not for semantic segmentation (Tab. 5(a)): removing 3D cross-view fusion layers results in a 6.9% mAP drop for instance segmentation, and a slight drop of 1.3% mIoU in semantic segmentation. Popular 2D-based 3D open vocabulary works [21, 37] without strong cross-view fusion only focus on semantic segmentation and thus could not uncover this issue. Row-3 shows a 5.7% mAP drop when cross-view 3D fusion happens after all within-view 2D layers instead of interleaving the within-view and cross-view fusion.
@@
2D pre-trained weight initialization helps (Tab. 5(b)): initializing only the image backbone with pre-trained weights, instead of all layers (except the 3D fusion layers), results in a 4.7% mAP drop (row-2). Starting the entire model from scratch leads to a larger drop of 9.1% mAP (row-3). This underscores the importance of sharing as many parameters as possible with the 2D models to leverage the maximum possible 2D pre-trained weights.
@@
Stronger 2D backbones helps (Tab. 5(c)): using Swin-B over ResNet-50 leads to significant performance gains, suggesting that ODIN can directly benefit from advancements in 2D computer vision.
@@
Finetuning everything including the pre-trained parameters helps (Tab. 5(c)): while ResNet50’s performance remains similar, Swin’s performance increases substantially when we fine-tune all parameters. Intuitively, unfreezing the backbone allows 2D layers to adapt to cross-view fused features better. Thus, we keep our backbone unfrozen in all experiments.
@@
Supplying 2D features directly to 3D models does not help: Concatenating 2D features with XYZ+RGB as input to Mask3D yields 53.8% mAP performance, comparable to 53.3%footnotetext: ${}^{\dagger}$e do not use the expensive DB-SCAN post-processing of Mask3D, and hence it gets 53.3% mAP instead of 55.2% as reported by their paper of the baseline model with only XYZ+RGB as input.
@@ 4.6 Limitations
Our experiments reveal the following limitations for ODIN: Firstly, like other top-performing 3D models, it depends on accurate depth and camera poses. Inaccurate depth or camera poses causes a sharp decrease in performance (similar to other 3D models, like Mask3D). As our experiments show, our model shines over variants in simulation, where camera poses and depths are accurate. In our future work, we aim to explore unifying depth and camera pose estimation with semantic scene parsing, thus making 3D models more resilient to noise. Secondly, in this paper, we limited our scope to exploring the design of a unified architecture without scaling-up 3D learning by training on diverse 2D and 3D datasets jointly. We aim to explore this in future in order to achieve strong generalization to in-the-wild scenarios, akin to the current foundational 2D perception systems.
@@ Appendix A Experiments
ODIN-Swin-B (Ours) 74.0
@@
Sensor RGBD Point Cloud ODIN-Swin-B (Ours) 36.4
@@ A.2 Evaluation on S3DIS and Matterport3D
S3DIS: S3DIS comprises 6 building-scale scenes, typically divided into 5 for training and 1 for testing. The dataset provides raw RGB-D images, captured panorama images, and images rendered from the mesh obtained after reconstructing the original sensor data. Unlike Matterport3D, S3DIS do not provide undistorted raw images; thus, we use the provided rendered RGB-D images. Some rooms in S3DIS have misalignments between RGB-D images and point clouds, which we address by incorporating fixes from DeepViewAgg [40] and introducing our own adjustments. Despite these fixes, certain scenes still exhibit significantly low overlap between RGB-D images and the provided mesh-sampled point cloud. To mitigate this, we query images from other rooms and verify their overlap with the provided point cloud for a room. This partially helps in addressing the low overlap issue.
The official S3DIS benchmark evaluates 13 classes. Due to the dataset’s small size, some models pre-train on additional datasets like ScanNet, as seen in SoftGroup [49], and on Structured3D datasets [59], consisting of 21,835 rooms, as done by Swin3D-L [55]. Similar to Mask3D [44], we report results in both settings of training from scratch and starting from weights trained on ScanNet. We use the same baselines as discussed in Sec. 4.1.
Like ScanNet and ScanNet200, both S3DIS and Matterport3D undergo post-processing of collected RGB-D data to construct a mesh, from which a point cloud is sampled and labeled. Consequently, following the approach outlined in Sec. 4.1, we train both Mask3D [44] and our model using RGB-D sensor point cloud data and evaluate on the benchmark-provided point cloud. Additionally, we explore model variants by training and testing them on the mesh-sampled point cloud for comparative analysis.
@@
ODIN-ResNet50 (Ours) 20.9 54.2 11.2 27.2
ODIN-Swin-B (Ours) 21.6 60.3 13.0 33.1
@@
Sensor RGBD Point Cloud ODIN-ResNet50 (Ours) 54.6 67.2 22.0 29.4
ODIN-Swin-B (Ours) 58.4 70.7 28.2 37.4
@@
(c) Comparison on S3DIS Area5 for Instance Segmentation Task. (${}^{\dagger}$ = uses additional data)
@@
Sensor RGBD Point Cloud Mask3D [44] 40.7 54.6 64.2
Mask3D [44] ${}^{\dagger}$ 41.3 55.9 66.1
ODIN-ResNet50 (Ours) 37.2 49.4 62.1
ODIN-ResNet50 ${}^{\dagger}$ (Ours) 46.1 61.2 74.1
ODIN-Swin-B ${}^{\dagger}$ (Ours) 42.7 58.6 71.3
Mesh Sampled Point Cloud SoftGroup [49] ${}^{\dagger}$ 51.6 66.1 -
@@
(d) Comparison on S3DIS for Semantic Segmentation Task. (${}^{\dagger}$ = uses additional data)
@@
Sensor RGBD Point Cloud MVPNet [20] 62.4
@@
ODIN-ResNet50 (Ours) 59.6
ODIN-ResNet50 ${}^{\dagger}$ (Ours) 69.6
ODIN-Swin-B ${}^{\dagger}$ (Ours) 67.3
Rendered RGBD Point Cloud VMVF [26] 65.4
@@
Swin3D-L [55] ${}^{\dagger}$ 74.5
ODIN outperforms SOTA 3D models with sensor point cloud input and closely matches models using mesh-sampled point clouds on Matterport3D Instance Segmentation Benchmark (Tab. 7(a)): and surpassing them on the mAP25 metric. In a long-tail distribution setting of 160 classes, ODIN outperforms mesh-sampled point cloud models, particularly excelling in the mAP25 metric. These findings align with the results reported in Sec. 4.1 for the ScanNet and ScanNet200 datasets.
@@
On S3DIS Instance Segmentation Benchmark (Tab. 7(c)), in the setup where baseline Mask3D start from ScanNet pre-trained checkpoint, our model outperforms them in the sensor point cloud setup but obtains lower performance compared to mesh sampled point cloud methods and when compared on the setup where all models train from scratch.
On S3DIS Semantic Segmentation Benchmark (Tab. 7(d), ODIN trained with ScanNet weight initialization outperforms all sensor or rendered RGBD point cloud based methods, while achieving competitive performance on mesh sampled point cloud. When trained from scratch, it is much worse than other baselines. Given the limited dataset size of S3DIS with only 200 training scenes, we observe severe overfitting. As a consequence, ODIN with stronger backbones, such as Swin-B, underperforms compared to ODIN with ResNet50.
@@ A.3 ScanNet200 Detailed Results
Table 8: Detailed ScanNet200 results for Instance Segmentation
@@
Sensor RGBD point cloud Mask3D ${}^{\dagger}$ [44] 15.5 21.4 24.3 21.9 31.4 37.1 13.0 17.2 18.9 7.9 10.3 11.5
ODIN-ResNet50 (Ours) 26.0 37.6 43.8 34.5 50.9 61.6 23.8 34.5 38.8 18.4 25.5 28.7
ODIN-Swin-B (Ours) 30.0 43.0 51.0 35.2 52.3 65.2 28.7 39.5 46.7 25.3 35.8 39.5
@@ A.5 Inference Time
We assess the inference time of Mask3D and ODIN by averaging the forward pass time of each model across the entire validation set, utilizing a 40 GB VRAM A100. When fed the mesh-sampled point cloud directly, Mask3D achieves an inference time of 228ms. When provided with the sensor point cloud as input, the inference time increases to 864 ms. This could be attributed to the additional computational overhead involved in transferring features from the sensor point cloud to the mesh-sampled point cloud before generating the output segmentation mask. Additionally, the sensor point cloud is typically larger than the mesh-sampled point cloud, as it also includes regions that were discarded during the mesh generation post-processing. ODIN-SwinB, which operates over the sensor point cloud, has an inference time of 960ms.
@@ Appendix B Additional Implementation Details
Image Resolutions We use a resolution of $240\times 320$ for ScanNet, $480\times 640$ for ScanNet200, and $512\times 512$ for AI2Thor. In our AI2THOR experiments, we discovered that employing higher image resolutions enhances the detection of smaller objects, with no noticeable impact on the detection of larger ScanNet-like objects. This observation was confirmed in ScanNet, where we experimented with $480\times 640$ image resolutions and did not observe any discernible benefit.

### v2

@@ 4.4 Ablations and Variants
Joint 2D-3D training helps 3D perception We compare joint training of ODIN on sensor RGB-D point clouds from ScanNet and 2D RGB images from COCO to variants trained independently on 2D and 3D data, all initialized from pre-trained COCO weights. Since there are different classes in ScanNet and COCO, we use our open-vocabulary semantic class-decoding head instead of the vanilla closed-vocabulary head. Results in Tab. 4 show that joint training yields a 1.3% absolute improvement in 3D, and causes a similar drop in 2D. This experiment indicates that a single architecture can perform well on both 2D and 3D tasks, thus indicating that we may not need to design vastly different architectures in either domain. However, the drop in 2D performance indicates a potential for further improvements in the architecture design to retain the performance in the 2D domain. Nevertheless, this experiment highlights the benefits of joint training with 2D datasets for 3D segmentation in ODIN. Note that we do not jointly train on 2D and 3D datasets in any of our other experiments due to computational constraints.
@@
Cross-View fusion is crucial for instance segmentation but not for semantic segmentation (Tab. 5(a)): removing 3D cross-view fusion layers results in an 8.5% mAP drop for instance segmentation, without any significant effect in semantic segmentation. Popular 2D-based 3D open vocabulary works [21, 37] without strong cross-view fusion only focus on semantic segmentation and thus could not uncover this issue. Row-3 shows a 6.1% mAP drop when cross-view 3D fusion happens after all within-view 2D layers instead of interleaving the within-view and cross-view fusion.
@@
2D pre-trained weight initialization helps (Tab. 5(b)): initializing only the image backbone with pre-trained weights, instead of all layers (except the 3D fusion layers), results in a 5.5% mAP drop (row-2). Starting the entire model from scratch leads to a larger drop of 6.3% mAP (row-3). This underscores the importance of sharing as many parameters as possible with the 2D models to leverage the maximum possible 2D pre-trained weights.
@@
Stronger 2D backbones helps (Tab. 5(c)): using Swin-B over ResNet50 leads to significant performance gains, suggesting that ODIN can directly benefit from advancements in 2D computer vision.
@@
Finetuning everything including the pre-trained parameters helps (Tab. 5(c)): ResNet50’s and Swin’s performance increases substantially when we fine-tune all parameters. Intuitively, unfreezing the backbone allows 2D layers to adapt to cross-view fused features better. Thus, we keep our backbone unfrozen in all experiments.
@@
Supplying 2D features directly to 3D models does not help: Concatenating 2D features with XYZ+RGB as input to Mask3D yields 53.8% mAP performance, comparable to 53.3%footnotetext: We do not use the expensive DB-SCAN post-processing of Mask3D, and hence it gets 53.3% mAP instead of 55.2% as reported by their paper of the baseline model with only XYZ+RGB as input.
@@ 4.6 Limitations
Our experiments reveal the following limitations for ODIN: Firstly, like other top-performing 3D models, it depends on accurate depth and camera poses. Inaccurate depth or camera poses cause a sharp decrease in performance (similar to other 3D models, like Mask3D). In our future work, we aim to explore unifying depth and camera pose estimation with semantic scene parsing, thus making 3D models more resilient to noise. Secondly, in this paper, we limited our scope to exploring the design of a unified architecture without scaling up 3D learning by training on diverse 2D and 3D datasets jointly. We aim to explore this in future to achieve strong generalization to in-the-wild scenarios, akin to the current foundational 2D perception systems. Our results suggest a competition between 2D and 3D segmentation performance when training ODIN jointly on both modalities. Exploring ways to make 2D and 3D training more synergistic is a promising avenue for future work.
@@ Appendix A Experiments
ODIN-Swin-B (Ours) 74.4
@@
Sensor RGBD Point Cloud ODIN-Swin-B (Ours) 36.8
@@ A.2 Evaluation on S3DIS and Matterport3D
S3DIS: S3DIS comprises 6 building-scale scenes, typically divided into 5 for training and 1 for testing. The dataset provides raw RGB-D images, captured panorama images, and images rendered from the mesh obtained after reconstructing the original sensor data. Unlike Matterport3D, S3DIS do not provide undistorted raw images; thus, we use the provided rendered RGB-D images. Some rooms in S3DIS have major misalignments between RGB-D images and point clouds, which we partially address by incorporating fixes from DeepViewAgg [40] and introducing our own adjustments. Despite these fixes, certain scenes still exhibit significantly low overlap between RGB-D images and the provided mesh-sampled point cloud. To mitigate this, we query images from other rooms and verify their overlap with the provided point cloud for a room. This partially helps in addressing the low overlap issue.
The official S3DIS benchmark evaluates 13 classes. Due to the dataset’s small size, some models pre-train on additional datasets like ScanNet, as seen in SoftGroup [49], and on Structured3D datasets [59], consisting of 21,835 rooms, as done by Swin3D-L [55]. Similar to Mask3D [44], we report results in both settings of training from scratch and starting from weights trained on ScanNet.
Like ScanNet and ScanNet200, both S3DIS and Matterport3D undergo post-processing of collected RGB-D data to construct a mesh, from which a point cloud is sampled and labeled. Hence, we train both Mask3D [44] and our model using RGB-D sensor point cloud data and evaluate on the benchmark-provided point cloud. Additionally, we explore model variants by training and testing them on the mesh-sampled point cloud for comparative analysis.
@@
ODIN-ResNet50 (Ours) 22.5 56.4 11.5 27.6
ODIN-Swin-B (Ours) 24.7 63.8 14.5 36.8
@@
Sensor RGBD Point Cloud ODIN-ResNet50 (Ours) 54.5 65.8 22.4 28.5
ODIN-Swin-B (Ours) 57.3 69.4 28.6 38.2
@@
(c) Comparison on S3DIS Area5 for Instance Segmentation Task. ( = uses additional data)
@@
RGBD Point Cloud Mask3D [44] 40.7 54.6 64.2
Mask3D [44] 41.3 55.9 66.1
ODIN-ResNet50 (Ours) 36.3 48.0 61.2
ODIN-ResNet50 (Ours) 44.7 57.7 67.5
ODIN-Swin-B (Ours) 43.0 56.4 70.0
Mesh Sampled Point Cloud SoftGroup [49] 51.6 66.1 -
@@
(d) Comparison on S3DIS for Semantic Segmentation Task. ( = uses additional data)
@@
RGBD Point Cloud MVPNet [20] 62.4
VMVF [26] 65.4
@@
ODIN-ResNet50 (Ours) 59.7
ODIN-ResNet50 (Ours) 66.8
ODIN-Swin-B (Ours) 68.6
@@
Swin3D-L [55] 74.5
ODIN outperforms SOTA 3D models on Matterport3D Instance Segmentation Benchmark across all settings (Tab. 7(a))
@@
On S3DIS Instance Segmentation Benchmark (Tab. 7(c)), in the setup where baseline Mask3D start from ScanNet pre-trained checkpoint, our model outperforms them in the RGBD point cloud setup but obtains lower performance compared to mesh sampled point cloud methods and when compared on the setup where all models train from scratch.
On S3DIS Semantic Segmentation Benchmark (Tab. 7(d), ODIN trained with ScanNet weight initialization outperforms all RGBD point cloud based methods, while achieving competitive performance on mesh sampled point cloud. When trained from scratch, it is much worse than other baselines. Given the limited dataset size of S3DIS with only 200 training scenes, we observe severe overfitting.
@@ A.3 ScanNet200 Detailed Results
Table 8: Detailed ScanNet200 results for Instance Segmentation (§ = trained by us using official codebase)
@@
Sensor RGBD point cloud Mask3D § [44] 15.5 21.4 24.3 21.9 31.4 37.1 13.0 17.2 18.9 7.9 10.3 11.5
ODIN-ResNet50 (Ours) 25.6 36.9 43.8 34.8 51.1 63.9 23.4 33.4 37.9 17.8 24.9 28.1
ODIN-Swin-B (Ours) 31.5 45.3 53.1 37.5 54.2 66.1 31.6 43.9 50.2 24.1 36.6 41.2
@@ A.5 Inference Time
We assess the inference time of Mask3D and ODIN by averaging the forward pass time of each model across the entire validation set, utilizing a 40 GB VRAM A100. When fed the mesh-sampled point cloud directly, Mask3D achieves an inference time of 228ms. When provided with the sensor point cloud as input, the inference time increases to 864 ms. Mask3D with sensor point cloud is slower than with mesh point cloud because at the same voxel size (0.02m), more voxels are occupied in sensor point cloud ( 110k on avg.) compared to mesh point clouds ( 64k on avg.) as mesh-cleaning sometimes discards large portion of the scene. The transfer of features from the sensor point cloud to the mesh point cloud adds an extra 7 ms. ODIN-SwinB, which operates over the sensor point cloud, has an inference time of 960ms.
@@ Appendix B Additional Implementation Details
Image Resolutions We use a resolution of $256\times 256$ for ScanNet, $512\times 512$ for ScanNet200, and AI2THOR. In our AI2THOR experiments, we discovered that employing higher image resolutions enhances the detection of smaller objects, with no noticeable impact on the detection of larger ScanNet-like objects. This observation was confirmed in ScanNet, where we experimented with $512\times 512$ image resolutions and did not observe any discernible benefit.